OpenStudy (anonymous):

integrate t^2sqrt(9t^4+4t^2) does anybody know how i would go about doing this with u substitution?

OpenStudy (anonymous):

\[\int\limits_{0}^{1}t^2\sqrt{9t^4+4t^2}dt\]

OpenStudy (anonymous):

first off; you can factor the t^2 from the sqrt outside to get t

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

\[t^3\sqrt{9t^2+4}\]

OpenStudy (anonymous):

this you can rewrite as \[3t ^{3}\sqrt{t ^{2}+(\frac{ 2 }{ 3 })^{2}}\]

OpenStudy (anonymous):

now you can think of a trig substitution

OpenStudy (anonymous):

trig sub? theres gotta be a way to do this using u-sub. >.<

OpenStudy (anonymous):

\[t = \frac{ 2 }{ 3 }\tan \Theta \]

OpenStudy (anonymous):

use this substitution; trig identity will come in handy

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20101002183858AAV3Jqm i think theres an easier way to go about doing this. when i look at this though, i cant keep track of what the person is doing

OpenStudy (anonymous):

it will get reduced to \[\frac{ 32 }{ 81 }\int\limits \tan ^{4}\Theta \sec ^{2}\Theta d \Theta \]

OpenStudy (anonymous):

now make the substitution \[u = \tan \Theta \]

OpenStudy (anonymous):

you'll get \[\frac{ 32 }{ 81 }\int\limits u ^{4} du\]

OpenStudy (anonymous):

which you can now easily solve

OpenStudy (anonymous):

then plug back u in terms of tan theta

OpenStudy (anonymous):

and finally plug back in tan theta in terms of t

OpenStudy (anonymous):

and you get the final ans

OpenStudy (anonymous):

alrite thanks. are you 100% postive there isnt a way to do this using just u-sub?

OpenStudy (anonymous):

oh wait; we are doing definite integral, so whenever you make a substitution the limits on the integral change

OpenStudy (anonymous):

in this case it will change twice since we make two substitutions

OpenStudy (anonymous):

so in the final step just evaluate the integral to get a number